# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

### v2

@@ 3.1 A toy model background
In order to obtain the MZW-model from the transverse equations of motion at leading order, like (2.6), we specify $\mathcal{M}_{H}$ to be a Lie group $G$.
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}\kappa_{ab}e^{a}e^{b}$ (3.1)
@@
$e^{a}={e_{m}}^{a}\mathrm{d}\hat{X}^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}\hat{X}^{0}$ or $v=\mathrm{d}\hat{X}^{i}$, for $i=1,2$. This choice will be related to the choice of $v$ in the MZW-model (1.4). In the same way as in the previous section, at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$, are
$0$ $=\partial^{N}j_{N}^{a}-\frac{1}{2}{f^{a}}_{bc}j^{b}_{M}j^{c}_{N}v_{L}\epsilon^{MNL},$ (3.2)
for the expansion (2.3). We recognise the MZW-model’s equation of motion (1.4) for the transverse directions. The longitudinal equations of motion completely decouples at leading order as in (2.5).
@@ 3.2 Embedding into a supergravity solution
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{R}{\bar{R}}\sigma^{0},0,r_{0}\right)$ (3.4)
@@
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is cured by the parameterisation of $t(\sigma)$ including the ratio $\frac{R}{\bar{R}}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.7), one reproduces the MZW-model equations of motion for the fluctations $X_{S^{3}}^{m}$.
@@ 4 Outlook
In this letter, we proposed a construction that shows that a certain regime of the (bosonic) membrane can be described by a classically integrable theory, the $(1+2)$-dimensional MZW-model, which has been studied in great detail in the literature. In this way, this proposal also offers a physical realisation of this model. Some crucial questions for future research will be:
@@
Moreover, for the integrable two-dimensional $\sigma$-models, generalised geometry resp. O(d,d)(d,d)-duality, was used in order to construct and understand new integrable models [37, 38, 39]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}$-structure, has been obtained and discussed in the literature [40, 41, 42, 43, 44, 45]. It might be fruitful to systematically study non-Lorentzian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [30].
@@ Acknowledgements
The research of D.O. is part of project No. 2024/55/D/ST2/01205 funded by the National Science Centre (NCN) of Poland. The author thanks the Institute for Theoretical Physics of ETH Zurich for their warm hospitality when finalising this work, the participants of the ’Workshop on Higher-d Integrability’ in Favignana for fruitful discussions about the draft, and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346. Special credits belong to the anonymous referee of Physical Review Letters pointing out a crucial mistake in the original draft.
@@ Appendix A The non-Lorentzian expansion of supergravity
The non-Lorentzian nature of the MZW-model suggests the use of non-Lorentzian geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [46, 47] and gives rise to the following expansions of metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter:
@@
Typically, the non-relativistic limit $c\rightarrow\infty$ was considered in the 11d supergravity case, but in other cases also the hyper-relativistic or Carollian limit ($c\rightarrow 0$) was considered from similar expansions in general relavitiy [48]. Both cases are of interest in main text.

### v3

@@ 3.1 A toy model background
In order to obtain the MZW-model from the setting in the previous section, we specify $\mathcal{M}_{H}$ to be a Lie group $G$, in particular we choose:
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}\kappa_{ab}e^{a}e^{b},$ (3.1)
@@
$e^{a}={e_{m}}^{a}\mathrm{d}\hat{X}^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one-form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}\hat{X}^{y}$ for $y=0,1$ or $2$. This choice will be related to the choice of $v$ in the MZW-model (1.2). In particular, one is left with
$\int C_{3}=\frac{1}{3}\omega R\bar{R}^{2}\int\sigma^{y}\text{tr}\left(j\wedge[j\ \overset{\wedge}{,}\ j]\right)+\mathcal{O}(\omega^{0})$ (3.2)
on the world-volume after applying (2.3).
One can easily see that for this particular background the $S_{H}$, the part of the membrane action (2.8) containing the transverse fluctuations, becomes the MZW model action
$S_{H}=S_{MZW}.$ (3.3)
In appendix B, we present how a similar statement can be obtained for the Polyakov formulation of the membrane at the level of the equations of motion.
@@ 3.2 Embedding into a supergravity solution
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{R}{\bar{R}}\sigma^{0},0,0\right)$ (3.5)
@@
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions, but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is rectified by parameterisation of $t(\sigma)$ including the ratio $\frac{R}{\bar{R}}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.11), one reproduces the action of the MZW-model for fluctations $X_{S^{3}}^{m}$.
@@ 4 Outlook
In this letter, we proposed a construction that shows that a certain regime of the (bosonic) membrane can be described by a classically integrable theory, the $(1+2)$-dimensional MZW-model, which has been studied in some detail in the literature. In this way, this proposal also offers a physical realisation of this model. Some crucial questions for future research will be
@@
[…]
For integrable two-dimensional $\sigma$-models, generalised geometry resp. O(d,d)(d,d)-duality was used in order to construct and understand new integrable models [38, 39, 40]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}$-structure, has been obtained and discussed in the literature [41, 42, 43, 44, 45, 46]. It might be fruitful to systematically study non-Lorentzian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [30].
@@ Acknowledgements
The research of D.O. is part of project No. 2024/55/D/ST2/01205 funded by the National Science Centre (NCN) of Poland. The author thanks Arkady Tseytlin for comments on the first version, the Institute for Theoretical Physics of ETH Zurich for their hospitality when finalising this work, the participants of the ’Workshop on Higher-d Integrability’ in Favignana for fruitful discussions about the draft, and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346. Special credits belong to to the anonymous referees of Physical Review Letters for their crucial constructive comments on the original manuscript.
@@ Appendix A The non-Lorentzian expansion of supergravity
The non-Lorentzian nature of the MZW-model suggests the use of non-Lorentzian geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [47, 48] and gives rise to the following expansions of the metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter:
@@
For this expansion, only the non-relativistic limit $c\rightarrow\infty$ was considered in the literature. We take this as a heuristic motivation that an ultra-relativistic or Carollian limit ($c\rightarrow 0$) should take a similar form, motivated by similar considerations in general relavitiy [49]. Both non- and ultra-relativistic limits are of interest in main text.
@@
Appendix B The MZW-model from the Polyakov formulation
The standard action [2] of the bosonic membrane in the Polyakov form including a dynamical world-volume metric $\gamma$ is
$S$ $=T\int\mathrm{d}^{3}\sigma\left(\frac{1}{2}\sqrt{-\gamma}\gamma^{MN}\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}\right$. (B.1)
with tension $T$ and in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $\hat{X}^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric:
$\gamma_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}.$ (B.2)
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}\hat{X}^{\nu}\partial_{K}\hat{X}^{\kappa}\partial_{L}\hat{X}^{\lambda},$ (B.3)
where $\Gamma$ denote Christoffel symbols and $F=\mathrm{d}C$. Due to the non-trivial nature of the world-volume metric $\gamma$, these equations are quite intractable, in general. After restricting to the general toy model background (2.2) and wrapping ansatz (2.3) for $\omega\rightarrow\infty$ and $\frac{\bar{R}^{2}}{R^{2}\omega^{2}}\rightarrow 0$ the induced metric becomes (2.4).
Including this into the equations of motion (B.3) for $X^{M}$ $X^{m}$, one obtains at leading order
$0$ $=\omega R\left(0+\mathcal{O}\left(\frac{1}{\omega},\frac{\bar{R}^{2}}{R^{2}\omega}\right)\right)$ (B.4)
$0$ $=\omega R\left(\partial_{M}\partial^{M}X^{m}+\Gamma_{(H)}^{m}{}_{kl}\partial_{M}X^{k}\partial_{N}X^{l}\right$. (B.5)
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\frac{1}{\omega},\frac{\bar{R}^{2}}{R^{2}\omega}\right)\right).$
If one specifies to the special background of the form (3.1) ($\mathcal{M}_{H}$ being a Lie group), at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$ become the equations of motion of the MZW model (1.2).
Remarkably, at leading order the longitudinal equations of motion vanish identically – at higher order one would expect additional coupling of longitudinal and transversal fluctuations. But we should note that in this formulation the relative and absolute scaling of both equations is not obvious – as compared to the formulation in terms of the action (2.8).
As argued in the main text, the limit does not seem to commute with classical reformulations (such as reparameterisations), and also seemingly the formulation as Polyakov or Nambu-Goto type model. Nevertheless, the defining feature of this limit (reproducing the MZW model) is present in any case.